# Benchmarking the Spatial Robustness of DNNs via Natural and Adversarial Localized Corruptions

### v1

@@ 6.1 Experimental Setup
Metric $M$ $\overline{M}$ $M$ $\overline{M}$ $M$ $\overline{M}$ $M$ $\overline{M}$ $M$ $\overline{M}$ $M$ $\overline{M}$ $M$ $\overline{M}$ $M$ $\overline{M}$ $M$ $\overline{M}$
@@
Table 1: Analysis of natural corruptions considering brightness, snow, and gaussian noise as corruption types at different severity levels (1, 2, 3). For comparison, the first row reports the clean average accuracy across the entire image (MeanAcc score), while the remaining columns analyze the accuracy within and outside the corrupted region, represented as $M$ and $\overline{M}$, respectively, following the analysis presented in Section 3. The corruption ratio is fixed at 50% of the image, and the same seed is applied for each test to ensure fair comparisons. For each metric and analysis, the top-2 models are highlighted in green for easier interpretation.
@@
Figure 3: Illustrations of the effects of localized natural corruptions across different semantic segmentation models (BiseNetX39, PSPNet, SegFormer1, DeepLabV3-ResNet) for localized natural corruptions with a ratio of $30\%$ and severity level 3. Synthetic snow is shown in the third row, and Gaussian noise in the second row, while the clean image and corresponding outputs are displayed in the first row.
6.2 Evaluation of Natural Robustness
@@
We first evaluate the natural robustness of the selected SS models. Table 1 reports the results for the Non-Corrupted Region Analysis ($\overline{M}$ in the table) and the Corrupted Region Analysis ($M$ in the table), considering different localized natural corruptions (synthetic snow, brightness adjustments, and gaussian noise) across different severity levels. For these tests, the corruption ratio is fixed to $r=0.5$, and corruptions are applied by splitting the images into patches of size $(256,256)$. To ensure fairness in the evaluation process, the seed was consistently initialized for each test. To improve table readability, we highlighted the top-2 models for each column.
@@
Considering the robustness in the corrupted region (columns labeled with $M$), for snow and brightness adjustments, no model demonstrates a clear and consistent superiority over the others, suggesting that convolution-based architectures remain valid alternatives to SegFormers. This observation does not hold in the case of gaussian noise, where transformer-based architectures exhibit a significant accuracy advantage in the corrupted region. Regarding the robustness in the non-corrupted region (columns labeled with $\overline{M}$), transformer-like architectures achieve significnatly better performance, especially when addressing higher severity levels of localized corruptions. In these cases, natural corruptions have the potential to substantially affect non-corrupted regions, thereby reducing the accuracy of $A_{\overline{M}}$. However, SegFormer demonstrates robustness by preventing the spread of mispredictions to otherwise clean regions.
@@
In Figure 3, we also show the outputs of the models when localized natural perturbations are applied, while the clean image and corresponding outputs of different models are displayed in the first row. For this test, we use a corruption ratio of $r=0.3$ and a severity level of $3$. The corresponding results align with those in the table and previous observations, showing that, in general, the SegFormer architecture achieves better robustness both within and outside the corrupted region.
@@
Figure 4: Analysis of the impact of the corruption ratio. We report the relative error with respect to the non-corrupted case, evaluated within the corrupted region (left plots) and the non-corrupted region (right plots). In the top plots, we use synthetic snow with severity level 3, and in the bottom plots, we use Gaussian noise with the same severity level.
@@
To further analyze the impact of localized corruption, we report in Figure 4(a) and Figure 4(b) the variation of the RCE (computed with respect to the clean accuracy in the addressed region, as described in the experimental setup) as a function of the corruption ratio. The RCE was computed both within (left plots) and outside (right plots) the corrupted regions, considering synthetic snow and Gaussian noise, respectively.
As shown in the plots, the corruption error follows different trends across the models. For small corruption ratios and corruption types with limited impact, such as synthetic snow, convolution-based models such as DeepLab-MobileNet and DeepLab-ResNet achieve performance comparable to SegFormer, both within and outside the corrupted region. However, when the corruption ratio increases, the error rises more rapidly for the convolution-based models, while SegFormer demonstrates superior robustness. As discussed above, a different behavior is observed when considering gaussian noise, for which SegFormer already exhibits significantly higher robustness even for small corrupted regions. We believe that this advantage is largely attributed to the application of global attention mechanisms in transformers, which can effectively attend to uncorrupted areas by leveraging features from those regions or, when regions are completely perturbed, distribute robust features more efficiently across the image [13].
@@
Figure 5: Accuracy in the non-corrupted region with and without (light color) localized adversarial perturbations using $\epsilon=16/255$ and $\epsilon=32/255$. The patch is always applied at the center of the image, with a size of (100, 100) in (a) and (200, 200) in (b), respectively.
@@
Figure 7: Illustration of the impact of using the Region-aware multi-attack analysis (Alg.1) on Segformer (a) and PSPNet (b) for two different images in the Cityscapes validation set. The use of multiple attacks allows for coverage of different parts of the input pixels, highlighting the complexity of solving a multi-objective adversarial optimization problem for semantic segmentation (SS) and demonstrating how the proposed algorithm can overcome this challenge. The attacks shown here use patches of size (200, 200) with $\epsilon=32/255$.
@@ 6.3 Evaluation of Adversarial Robustness
To further understand localized robustness, we evaluate the performance of the semantic segmentation (SS) models using the proposed region-aware multi-attack analysis (Algorithm 1), aiming for the identification of the worst-case scenario for spatial robustness.
@@
As discussed in Section 5, for adversarial evaluation, it is important to note that we focus on the analysis of non-corrupted regions ($A_{\overline{M}}$) only. This is because evaluating corrupted regions under localized attacks is straightforward, as any model will typically exhibit near-zero accuracy within the corrupted area. In contrast, analyzing the non-corrupted regions provides a more meaningful assessment of spatial robustness.
@@
In Figure 5(a) and Figure 5(b), we report the accuracy in the non-corrupted regions under attack, compared to the non-attacked case (‘clean’, shown in the lightest color in the bars), using localized adversarial attacks with patch sizes of (100, 100) and (200, 200), centered in the image, respectively. In both cases, we tested attacks with $\epsilon$ values of $16/255$ and $32/255$ (indicated by different colors in the plots). The multi-attack of Algorithm 1 is performed with $N_{\text{Att}}=3$ attacks and 50 optimization iterations (see Eq. (10)) per attack. Further analysis of these parameters is discussed in the following.
As shown in the plots, the trend is completely opposite to what we observed for natural corruptions. In this case, transformers exhibit lower robustness, primarily due to the global attention mechanisms they employ [13, 34, 33]. For example, even with a small patch size (100, 100), the accuracy drops significantly from 0.73 to 0.25 and 0.17 for $\epsilon=16/255$ and $\epsilon=32/255$, respectively.
@@
In contrast, convolution-based models, such as DeepLabV3 with ResNet backbones, demonstrate significantly higher robustness. This robustness can be attributed to the limited local receptive field inherent to convolutional layers [21]. However, it is important to note that these models are not purely convolutional, as they also integrate attention mechanisms to enhance performance. This integration can result in substantial accuracy drops, as seen with PSPNet when using patches of size (200, 200). This observation is consistent with previous studies, which suggest that even convolutional models employing high-context exploitation through attention-based mechanisms (such as pyramidal squeezing and excitation blocks) can still be vulnerable to certain localized adversarial perturbations [29, 24].
@@
Figure 8: Analysis of the impact of the proposed localized adversarial analysis when considering multiple attacks $N_{\textit{Att}}$ in Algorithm 1(a), with the number of iterations fixed to 50 and $\epsilon=16/255$ using patches of size (200, 200); and varying the number of optimization iterations per attack in (b), with the number of attacks fixed to 3 and the same settings for $\epsilon$ and patch size.
@@
Specifically, with a single attack (#1), the multi-attack is equivalent to running a standard PGD attack over the entire targeted region. However, as the number of attacks increases, the algorithm begins focusing on fooling regions that have not yet been attacked, thereby expanding the coverage to new regions that were previously unaffected. In fact, as one may expect, the effectiveness of the multi-attack significantly improves with more attacks, as indicated by the increase in RCE. This effect is particularly pronounced in models with high attention mechanisms, such as transformers and PSPNet, demonstrating the importance of the multi-attack approach in identifying worst-case scenarios of localized adversarial perturbations. For example, the RCE of SegFormer models increases significantly, rising from approximately 0.6 with a single attack to around 0.95 when using five attacks ($N_{\text{Att}}=5$).
We also provide illustrations of the effect of multiple attacks for SegFormer and PSPNet in Figures 7(a) and 7(b). As shown in the plots, the adoption of multiple attacks with awareness of the fooling regions allows for overcoming issues in the multi-objective adversarial problem for semantic segmentation (SS). In Figure 7(a), at each attack, the untargeted multi-attacks target different classes, highlighting the inter-class problem discussed previously in Figure 2(a). Similarly, for PSPNet (Figure 7(b)), addressing different fooling regions leads to a third attack capable of affecting the entire image, thereby overcoming potential objectives that previously limited the effectiveness of earlier attacks.
@@
We also analyze in Figure 8(b) the impact of varying the number of optimization iterations used at line 9 in Algorithm 1, fixing the number of attacks $N_{\text{Att}}=3$. As expected, the RCE increases when using more optimization iterations, such as 200 or 250, highlighting that the complexity of the multi-objective optimization problem in this context requires a higher number of iterations to improve effectiveness. However, we acknowledge that for large-resolution datasets, as Cityscapes, a limited number of iterations (e.g., 50) is sufficient to provide meaningful and consistent comparisons, as demonstrated in the plots. This balance helps find a trade-off between computational cost and robustness evaluation, particularly when dealing with high-resolution images, which are computationally expensive due to the gradient computations required to update the attack at each step.
@@
Finally, we highlight the importance of the position of the adversarial patch within the image. The previous attacks were conducted with the patch placed at the center of the image, which represents the worst-case position for standard convolutional models due to its ability to affect the largest possible receptive field [21, 28, 24]. However, it is important to analyze how the effectiveness of localized adversarial attacks decreases when the patch is shifted to more constrained areas, such as the corners of the image.
@@
Figure 11: Analysis of the natural and adversarial errors, computed as discussed in Equations 12 and 13, respectively, and clean accuracy when considering different models as $f_{1}$ and $f_{2}$ in the ensemble strategy defined in Equation 11, for different configurations of $\gamma$. The results aim to explore the existence of a possible trade-off to balance robustness against localized natural corruptions and adversarial perturbations.
@@ 6.4 Searching for a Trade-off in Ensembling
Given the results obtained in the above analyses of natural perturbations and adversarial attacks, we acknowledge that transformer architectures generally exhibit greater robustness to localized natural corruptions by effectively reducing misclassifications, not only within the corrupted regions but also in the surrounding non-corrupted areas. However, their reliance on high attention mechanisms makes them more vulnerable to localized adversarial attacks, resulting in a significant drop in accuracy compared to convolutional models.
Following this observation, we emphasize the importance of searching for a trade-off that balances natural robustness and worst-case adversarial robustness, which could be particularly beneficial in safety-critical domains such as autonomous driving. To address this challenge, we investigate the use of model ensembling, with natural and adversarial localized robustness as key metrics to evaluate the effectiveness of the ensemble strategy. To explore this approach, we first define a test-time weighted ensemble strategy [12] as:
@@
Specifically, for our tests, we analyze pairs of models where $f_{1}$ is a model with higher robustness to localized adversarial attacks (e.g., DeepLab-ResNet101, DeepLab-MobileNet, BiSeNet-ResNet101, and DDRNet23), and $f_{2}$ is demonstrates strong robustness to natural corruptions but lower robustness to localized adversarial attacks (e.g., SegFormer_b0 or SegFormer_b1).
To understand the benefits of ensembling, we redefine the RCE for natural and adversarial perturbations, as follows, to account for the combination of $f_{1}$ and $f_{2}$. Most important to us are the RCE for natural perturbations within the corrupted area, i.e.,
@@
where $\mathcal{P}_{i}(M)=M_{i}/|M|$ and $C$ is a set of naural corruptions, and the RCE for adversarial perturbations outside the corrupted area, i.e.,
@@
where $\mathcal{P}_{i}(M)=\overline{M}_{i}/|\overline{M}|$ and $C$ is a set of adversarial corruptions.
The rationale behind these error definitions is that, since $f_{2}$ and $f_{1}$ are chosen as the best-performing models for natural and adversarial robustness, respectively, this formulation allows us to understand how varying $\gamma$ influences improvements in natural robustness and the corresponding drop in adversarial robustness.
The results of Figure 11 confirm the expected trend. As we increase $\gamma$, the ensemble model becomes more robust to natural perturbations but also increasingly susceptible to localized adversarial attacks. Note that, in this case, the attacks are performed directly on the ensemble model, considering the corresponding $\gamma$ value. Interestingly, for certain values (e.g., $\gamma=0.4$), the adversarial robustness is largely preserved, while we observe a significant reduction in the natural error. This phenomenon is particularly noticeable when DeepLab architectures are used as $f_{1}$, as the natural error decreases slightly before the adversarial error curve begins to rise. This demonstrates that a positive trade-off between natural and adversarial robustness can be found.

### v3

@@ 6.2 Evaluation of Natural Robustness
Metric MAcc $A_{M}$ $A_{\overline{M}}$ $A_{M}$ $A_{\overline{M}}$ $A_{M}$ $A_{\overline{M}}$ $A_{M}$ $A_{\overline{M}}$ $A_{M}$ $A_{\overline{M}}$ $A_{M}$ $A_{\overline{M}}$ $A_{M}$ $A_{\overline{M}}$ $A_{M}$ $A_{\overline{M}}$ $A_{M}$ $A_{\overline{M}}$
@@
Table 1: Results of the analysis for natural corruptions considering brightness, snow, and gaussian noise as corruption types at different severity levels (1, 2, 3). The first column reports the clean average accuracy (MAcc), while the remaining columns analyze the accuracy within and outside the corrupted region, represented as $A_{M}$ and $A_{\overline{M}}$, respectively, following the analysis presented in Section 3. The corruption ratio is fixed at 50% of the image. For each metric and analysis, the top-2 models are highlighted in green for easier interpretation.
@@
Figure 3: Illustrations of the effects of localized natural corruptions across different semantic segmentation models (BiseNetX39, PSPNet, SegFormer-B1, DeepLabV3-ResNet) with a perturbation ratio of $r=0.3$ and severity level 3. The first row shows the clean image along with its ground truth label and corresponding model outputs. The second row presents the same image after applying localized Gaussian noise, and the third row shows it with synthetic snow applied.
@@
We first evaluate the natural robustness of the selected SS models. Table 1 reports the results for the Non-Corrupted Region Analysis ($A_{\overline{M}}$ in the table) and the Corrupted Region Analysis ($A_{M}$ in the table), considering different localized natural corruptions (synthetic snow, brightness corruptions, and gaussian noise) across different severity levels. For these tests, the corruption ratio is fixed to $r=0.5$, and corruptions are applied by splitting the images into patches of size $(256,256)$.
@@
Considering the robustness in the corrupted region (columns labeled with $A_{M}$), for snow and brightness adjustments, no model demonstrates a clear and consistent superiority over the others, suggesting that convolution-based architectures remain valid alternatives to SegFormers. This observation does not hold in the case of gaussian noise, where transformer-based architectures exhibit a significant accuracy advantage in the corrupted region.
Regarding the robustness in the non-corrupted region (columns labeled with $A_{\overline{M}}$), transformer-like architectures achieve significantly better performance, especially when addressing higher severity levels of localized corruptions. In these cases, natural corruptions have the potential to substantially affect non-corrupted regions, thereby reducing the accuracy of $A_{\overline{M}}$. However, SegFormer demonstrates robustness by preventing the spread of mispredictions to otherwise clean regions.
@@
In Figure 3, we also show the outputs of the models when localized natural perturbations are applied, while the clean image and corresponding outputs of different models are displayed in the first row. For this test, we use a corruption ratio of $r=0.3$ and a severity level of $3$. The corresponding results align with those in Table 1 and previous observations, showing that, in general, the SegFormer architectures achieve better robustness both within and outside the corrupted regions.
@@
Figure 4: Analysis of the impact of the corruption ratio. We report the relative corruption error RCE with respect to the corrupted region (left plots) and the non-corrupted region (right plots). We used synthetic snow and gaussian noise, in the top and bottom plots, respectively, with severity level 3.
@@
To further analyze the impact of localized corruption, we report in Figure 4(a) and Figure 4(b) the variation of the RCE (computed with respect to the clean accuracy in the addressed region) as a function of the corruption ratio. The RCE was computed both within (left plots) and outside (right plots) the corrupted regions, considering synthetic snow and gaussian noise, respectively.
As shown in the plots, the corruption error follows different trends across the models. For small corruption ratios and corruption types with limited impact, such as synthetic snow, convolution-based models such as DeepLab-MobileNet and DeepLab-ResNet achieve performance comparable to SegFormer, both within and outside the corrupted region. However, when the corruption ratio increases, the error rises more rapidly for the convolution-based models, while SegFormer demonstrates superior robustness. As discussed above, a different behavior is observed when considering gaussian noise, for which SegFormer already exhibits significantly higher robustness even for small corrupted regions. We believe that this advantage is largely attributed to the application of global attention in transformers, which can effectively attend to uncorrupted areas by leveraging features from those regions or, when regions are completely perturbed, distribute robust features more efficiently across the image [gu2022vision].
@@
Figure 5: Analysis of the accuracy in the non-corrupted region with (dark color) and without (light colors) localized adversarial perturbations. Attacks were performed using $\epsilon=16/255$ and $\epsilon=32/255$. The patch is always applied at the center of the image, with a size of (100, 100) in (a) and (200, 200) in (b), respectively.
@@
Figure 7: Illustration of the impact of using the Region-aware multi-attack adversarial analysis (Algorithm 1) on Segformer (a) and PSPNet (b) for two different images in the Cityscapes validation set. The use of multiple attacks allows for coverage of different parts of the input pixels, highlighting the complexity of solving a multi-objective adversarial optimization problem for SS and demonstrating how the proposed algorithm can overcome this challenge. The attacks shown here use patches of size (200, 200) with $\epsilon=32/255$.
@@ 6.3 Evaluation of Adversarial Robustness
To further understand localized robustness, we evaluate the performance of the SS models using the proposed region-aware multi-attack adversarial analysis (Algorithm 1), aiming for the identification of the worst-case scenario for spatial robustness.
@@
As discussed in Section 5, for adversarial evaluation, it is important to note that we focus on the analysis of non-corrupted regions ($A_{\overline{M}}$). This is because evaluating corrupted regions under localized attacks is straightforward, as any model will typically exhibit near-zero accuracy within the corrupted area. In contrast, analyzing the non-corrupted regions provides a more meaningful assessment of spatial robustness.
Please note that, for completeness, concerning the adversarial analysis, the full list of parameters involved is: the patch size, the attack magnitude $\epsilon$, the number of attacks $N_{\text{Att}}$, the position of the attack (which is always centered unless explicitly stated otherwise), and the number of iterations for each attack.
@@
In Figure 5(a) and Figure 5(b), we report the accuracy in the non-corrupted regions under attack, compared to the non-attacked case (‘clean’, shown in the lightest color in the bars), using localized adversarial attacks with patch sizes of (100, 100) and (200, 200), centered in the image, respectively. In both cases, we tested attacks with $\epsilon$ values of $16/255$ and $32/255$ (indicated by different colors in the plots). The multi-attack of Algorithm 1 is performed with $N_{\text{Att}}=3$ attacks and 50 optimization iterations (see Equation 10) per attack. Further analysis of these parameters is discussed in the following.
@@
As shown in the plots, the trend is completely opposite to what we observed for natural corruptions. In this case, transformers exhibit lower robustness, primarily due to the global attention mechanisms they employ [gu2022vision, xie2021segformer, vaswani_attention_2017]. For example, even with a small patch size (100, 100), the accuracy drops significantly from 0.73 to 0.25 and 0.17 for $\epsilon=16/255$ and $\epsilon=32/255$, respectively. In contrast, convolution-based models, such as DeepLabV3 with ResNet backbone, demonstrate significantly higher robustness. This robustness can be attributed to the limited local receptive field inherent to convolutional layers [luo2016understanding]. However, it is important to note that these models are not purely convolutional, as they also integrate attention mechanisms to enhance performance. This integration can result in substantial accuracy drops, as seen with PSPNet when using patches of size (200, 200). This observation is consistent with previous studies, which suggest that even convolutional models employing high-context exploitation through attention-based mechanisms (such as pyramidal squeezing and excitation blocks) can still be vulnerable to certain localized adversarial perturbations [rossolini_tnnls_2023, nakka_indirect_2020].
@@
Figure 8: Analysis of the impact of the proposed localized adversarial framework under different attack numbers $N_{\textit{Att}}$ in Algorithm 1 (a), and under different number of iterations (b). In (a), the number of iterations is fixed to 50, while in (b), the number of attacks is fixed to 3. In both analyses, patches are of fixed size (200, 200) and $\epsilon=16/255$.
@@
Specifically, with a single attack (#1), the multi-attack is equivalent to running a standard attack over the entire targeted region. However, as the number of attacks increases, the algorithm begins focusing on fooling regions that have not yet been attacked, thereby expanding the coverage to new regions that were previously unaffected. In fact, as one may expect, the effectiveness of the multi-attack significantly improves with more attacks, as indicated by the increase in RCE. This effect is particularly pronounced in models with high attention mechanisms, such as transformers and PSPNet, demonstrating the importance of the multi-attack approach in identifying worst-case scenarios of localized adversarial perturbations. For example, the RCE of SegFormer models increases significantly, rising from approximately 0.6 with a single attack to around 0.95 when using five attacks ($N_{\text{Att}}=5$).
We also provide illustrations of the effect of multiple attacks for SegFormer and PSPNet in Figures 7(a) and 7(b). As shown in the plots, the adoption of multiple attacks with awareness of the fooling regions allows for overcoming issues in the multi-objective adversarial problem for SS. In Figure 7(a), at each attack, the untargeted multi-attacks target different classes, highlighting the inter-class problem discussed previously in Figure 2(a). Similarly, for PSPNet (Figure 7(b)), addressing different fooling regions leads to a third attack capable of affecting the entire image, thereby overcoming potential objectives that previously limited the effectiveness of earlier attacks.
@@
We also analyze in Figure 8(b) the impact of varying the number of optimization iterations used at line 9 in Algorithm 1, fixing the number of attacks $N_{\text{Att}}=3$. As expected, the RCE increases when using more optimization iterations, such as 200 or 250, highlighting that the complexity of the multi-objective optimization problem in this context requires a higher number of iterations to improve effectiveness. However, we acknowledge that for large-resolution datasets, as Cityscapes, a limited number of iterations (e.g., 50) is sufficient to provide meaningful and consistent comparisons, as demonstrated in the plots. This balance helps find a trade-off between computational cost and robustness evaluation, particularly when dealing with high-resolution images, which are computationally and memory expensive due to the gradient computations required to update the attack at each step.
@@
Finally, we highlight the importance of the position of the adversarial perturbed patch within the image. The previous attacks were conducted with the patch placed at the center of the image, which represents the worst-case position for standard convolutional models due to its ability to affect the largest possible receptive field [luo2016understanding, rossolini_iccps24, nakka_indirect_2020]. However, it is important to analyze how the effectiveness of localized adversarial attacks decreases when the patch is shifted to more constrained areas, such as the corners of the image.
@@
Figure 11: Analysis of the natural and adversarial errors, computed as discussed in Equations 13 and 14, respectively, and clean accuracy when considering different models as $f_{1}$ and $f_{2}$ in the ensemble strategy defined in Equation 12, for different configurations of $\gamma$. The results aim to explore the existence of a possible trade-off to balance robustness against localized natural corruptions and adversarial perturbations.
@@ 6.4 Searching for a Trade-off in Ensembling
Given the results obtained in the above analyses of natural perturbations and adversarial attacks, we acknowledge that the selected transformer architectures generally exhibit greater robustness to localized natural corruptions by effectively reducing misclassifications, not only within the corrupted regions but also in the surrounding non-corrupted areas. However, their reliance on high attention mechanisms makes them more vulnerable to localized adversarial attacks, resulting in a significant drop in accuracy compared to convolutional models tested.
Following this observation, we emphasize the importance of searching for a trade-off that balances natural robustness and worst-case adversarial robustness, which could be particularly beneficial in safety-critical domains such as autonomous driving. To address this challenge, we investigate the use of model ensembling, with natural and adversarial localized robustness as key metrics to evaluate the effectiveness of the ensemble strategy. To explore this approach, we first define a test-time weighted ensemble strategy [ensemble_GANAIE2022105151] as:
@@
Specifically, for our tests, we analyze pairs of models where $f_{1}$ is a model with higher robustness to localized adversarial attacks (e.g., DeepLab-ResNet101, DeepLab-MobileNet, BiSeNet-ResNet101, and DDRNet23), and $f_{2}$ demonstrates stronger robustness to natural corruptions but lower robustness to localized adversarial attacks (e.g., SegFormer_b0 or SegFormer_b1).
To understand the benefits of ensembling, we redefine the RCE for natural and adversarial perturbations, as follows, to account for the combination of $f_{1}$ and $f_{2}$. In particular, most important to us for this analysis, we used as metrics: (i) the RCE (Equation 11) for natural perturbations within the corrupted area, i.e.,
@@
where $\mathcal{P}_{i}(M)=M_{i}/\mathcal{S}$ and $C$ is a set of natural corruptions (we considered gaussian noise, with $r=0.5$); and (ii) the RCE for adversarial perturbations outside the corrupted area, i.e.,
@@
where $\mathcal{P}_{i}(M)=(1-M_{i})/\overline{\mathcal{S}}$ and $C$ is a localized adversarial corruption (we used the attack proposed in Algorithm 1, with $N_{\textit{Att}}=5,\epsilon=16/255$).
The rationale behind these error definitions is that, since $f_{1}$ and $f_{2}$ are chosen as the best-performing models for natural and adversarial robustness, respectively, this formulation allows us to understand how varying $\gamma$ influences improvements in natural robustness and the corresponding drop in adversarial robustness.
The results of Figure 11 confirm the expected trend. As we increase $\gamma$, the ensemble model becomes more robust to natural perturbations but also increasingly susceptible to localized adversarial attacks. Note that, in this case, the attacks are performed directly on the ensemble model to provide a fair evaluation, considering the corresponding $\gamma$ value. Interestingly, for certain values (e.g., $\gamma=0.4$), the adversarial robustness is largely preserved, while we observe a significant reduction in the natural error. This phenomenon is particularly noticeable when DeepLab architectures are used as $f_{1}$, as the natural error decreases slightly before the adversarial error curve begins to rise. This demonstrates that a positive trade-off between natural and adversarial robustness can be found.